at the end of the race, a runner decelerates from a velocity of 9m/s^2. how far does she travel in next 5s?




1
Expert's answer
2021-03-14T19:20:16-0400

I assume, the initial velocity was "v_0 = 9m\/s". Then the acceleration was (by definition):


"a = \dfrac{v_f-v_0}{t} = -\dfrac{v_0}{t}"

where "v_f = 0", is the final speed of the runner, "t = 5s" is the time she decelerated.

The distance travelled in this time was:


"d = v_0t - \dfrac{at^2}{2} = v_0t - \dfrac{v_0t}{2} = \dfrac{v_0t}{2}"

Thus, obtain:


"d = \dfrac{9m/s\cdot 5s}{2} = 22.5m"

Answer. 22.5 m.